# Ax-Prover: A Deep Reasoning Agentic Framework for Theorem Proving in Mathematics and Quantum Physics

### v2

@@ Abstract
We present Ax-Prover, a multi-agent system for automated theorem proving in Lean that can solve problems across diverse scientific domains and operate either autonomously or collaboratively with human experts. To achieve this, Ax-Prover approaches scientific problem solving through formal proof generation, a process that demands both creative reasoning and strict syntactic rigor. Ax-Prover meets this challenge by equipping Large Language Models (LLMs), which provide knowledge and reasoning, with Lean tools via the Model Context Protocol (MCP), which ensure formal correctness. To evaluate its performance as an autonomous prover, we benchmark our approach against frontier LLMs and specialized prover models on two public math benchmarks and on two Lean benchmarks we introduce in the fields of abstract algebra and quantum theory. On public datasets, Ax-Prover is competitive with state-of-the-art provers, while it largely outperform them on the new benchmarks . This shows that, unlike specialized systems that struggle to generalize, our tool-based agentic theorem prover approach offers a generalizable methodology for formal verification across diverse scientific domains. Furthermore, we demonstrate Ax-Prover’s assistant capabilities in a practical use case, showing how it enabled an expert mathematician to formalize the proof of a complex cryptography theorem.
@@ 1 Introduction
Developing Large Language Models (LLMs) that can reason reliably across scientific domains remains a central challenge for AI, both in academia and in industry. LLM-based formal reasoning systems have mainly been developed for mathematics, where they have achieved outstanding results [16, 13]. Recently, considerable effort has been put into training reasoning LLMs for formal theorem proving using Lean [19], an open-source programming language and interactive proof assistant. Together with its community-driven Mathlib library [33], Lean provides a rigorous setting where AI systems must engage with symbolic reasoning and structured formalization, building on an evolving body of mathematical knowledge. LLM provers such as the DeepSeek-Prover series [60, 61, 49], Kimina-Prover-72B [58], Goedel-Prover [35, 36], and Seed-Prover [13] have shown that specialized prover models can be distilled from frontier LLMs and trained for theorem proving in Lean, reaching state-of-the-art performance on math benchmarks like miniF2F [65] and PutnamBench [55]. Despite these results, these models face some key limitations. First, since they were mainly trained and tested in the the domain of mathematics, their ability to generalize beyond this domain remains unclear. Relatedly, they are usually trained on fixed versions of the fast-evolving Mathlib library, making them brittle to changes introduced in new Mathlib versions – such as new or renamed definitions – unless continuously re-trained which adds significant cost. [^1] For example, Deepseek-Prover-V2-671B was released on April 30, 2025 and, in our experiments, we observed that, for example, it uses the lemma sqrt_eq_iff_sq_eq, which was deprecated in favor of sqrt_eq_iff_eq_sq on March 3, 2025. Second, while training improves their ability to produce Lean proofs, it narrows their capabilities relative to general LLMs as they become unable to use external tools and engage in human–AI collaboration. Finally, it is hard to run them as they require high-spec machines and expertise to be successfully deployed and used. Together, these issues suggest that scaling increasingly large specialized provers may yield diminishing returns in both flexibility and usability.
@@
In contrast, general-purpose LLMs like Claude [5] and GPT [45] encode substantial knowledge across diverse domains (e.g., mathematics, physics, and computer science), while also exhibiting strong natural language understanding, problem solving skills, and interaction capabilities. Moreover, they are easily accessible through APIs, making them convenient to deploy and integrate into any workflow. Yet, they are not explicitly trained to formalize statements or construct proofs in Lean, nor can they natively interact with the Lean environment. This creates a sharp division: specialized provers are tightly integrated with Lean but narrow in scope and hard to use, whereas general-purpose LLMs are broad in scope and easily accessible but lack the ability to interface with the formal reasoning infrastructure required for theorem proving.
@@
To address this gap, we introduce Ax-Prover,222“Ax” stands for “axiomatic”, reflecting the base principles in mathematics and physics, the domains explored in this work. a new agentic workflow for theorem proving in Lean that leverages the Model Context Protocol (MCP) [41] to equip general-purpose LLMs with Lean tools from the lean-lsp-mcp repository [22]. Ax-Prover combines the reasoning capabilities of LLMs with the formal verification power of Lean. The LLM analyzes unproven theorems, proposes proof sketches, and generates step-by-step Lean code, while the Lean tools allow the LLM to inspect goals, search for relevant results, locate errors, and verify proofs – capabilities essential for rigorous formal theorem proving.
@@
Ax-Prover overcomes the main limitations of state-of-the-art provers. First, using frontier LLMs prevents domain overspecialization while the MCP interface lets the system work with any recent version of Mathlib without needing to be retrained. Second, it preserves tool-use and conversational abilities, enabling interactive collaboration with humans. Third, by leveraging existing frontier models, it sidesteps the need to host or deploy specialized systems.
@@
We evaluated Ax-Prover on two public datasets that include mathematics competition problems (NuminaMath-LEAN [44] and PutnamBench [55]) and introduce two new datasets to enable evaluation in new domains. The first one, AbstractAlgebra, focuses on algebraic structures such as groups, rings, and fields, testing the provers’ abilities to reason in a more abstract, research-oriented setting rather than the competition-driven style of existing datasets. The second one is QuantumTheorems, which represents an initial step toward automated theorem proving in a scientific domain beyond pure mathematics, evaluating the models’ formal reasoning in quantum mechanics. Our results show that Ax-Prover has competitive performance on PutnamBench and outperforms general-purpose LLMs not equipped with Lean tools as well as state-of-the-art specialized provers on the other datasets, with a significant margin on the ones we introduce. This gap underscores the potential of Ax-Prover to act as the key AI verification tool for mathematically grounded scientific reasoning.
@@
Our contributions are threefold: (i) We design Ax-Prover, a lightweight agentic workflow that connects general-purpose LLMs to Lean tools via MCP, and demonstrate that it performs on par with or surpasses both general-purpose LLMs and specialized provers across several scientific domains; (ii) We introduce new formalized Lean datasets covering abstract algebra and quantum physics, complementing existing benchmarks; (iii) We showcase Ax-Prover’s capabilities as an assistant through a use case where the system collaborated with an expert mathematician to formally verify the proof of a recent cryptography result.
@@ 2 Related Work
Automated theorem proving in Lean has roots in classical approaches such as decision procedures [18, 10] and heuristic-guided proof search [30, 51], but they face challenges, specifically the former does not handle general mathematical domains (e.g. transcendental functions and complex numbers) and the latter do not before well out of distribution. More recent work integrates machine learning: from heuristic tuning [56] to premise selection and tactic prediction [28, 27], culminating in transformer-based language models capable of generating Lean proofs [48, 32, 47, 62]. More recent large-scale systems extend this trend by training LLMs on formal proof though distillation, supervised finetuning and reinforcement learning. Current examples of specialized models are Kimina-Prover [58], the DeepSeek-Prover family [60, 61, 49], Goedel-Prover 1 and 2 [35, 36], Prover Agent [9], Apollo [46], and Seed-Prover [13]. All of these are highly specialized provers, which take input a Lean theorem and autonomously produce a proof. A related line of work is exploring agentic flows that include frontier LLMs and a formal verifier; example include Hilbert [57] and Aristotle [2]. Although we also adopt a similar approach, some crucial differences exist, namely: (i) we give the LLM direct access to Lean tools via MCP; (ii) our framework requires no training [2] and does not rely on any specialized provers [57]; (iii) we demonstrate the effectiveness of our approach on domains other than mathematics; (iv) we showcase the capabilities of our system as an interactive assistant to human researchers.
Finally, a parallel line of work has explored classical machine learning for supporting experts in theorem proving in Lean, for example, in premise selection and tactic prediction [25, 11], and more recently through LLMs that connect to Lean via external interfaces [7, 8, 53]. These approaches illustrate the promise of AI-assisted proving, but they remain resource-intensive and difficult to adapt across scientific domains. Recent efforts, such as [31], attempt to remedy this by emphasizing greater adaptability within Lean. At the same time, there is growing interest in human–AI collaboration: conversational assistants [17] and “copilot”-style integrations [14] suggest how formal tools can augment, rather than replace, human reasoning. Our work builds on this trajectory by closing the gap between heavyweight, specialized provers and lightweight, researcher-friendly systems that can be more readily adapted to the evolving Lean ecosystem.
@@ 3 System Architecture
We implement Ax-Prover as a multi-agent architecture with three agents, each of which is implemented as an LLM with specific prompts: the Orchestrator, the Prover, and the Verifier. Following recent agentic designs for complex tasks such as scientific discovery [26, 63], we avoid a monolithic setup by assigning each specialized agent a distinct role. This separation enables specialization and modularity: agents can be independently optimized, replaced, or extended, allowing researchers to adapt Ax-Prover to their own needs without destabilizing the system.
@@ 3.1.1 Orchestrator
The Orchestrator’s role is analogous to a scheduler in distributed systems: it does not perform computation itself but ensures that computation flows smoothly across agents. It holds three main responsibilities. First, it handles task assignment, as it receives user input and instructs the Prover accordingly. Next, it manages feedback routing by taking diagnostic outputs from the Verifier and giving structured feedback to the Prover (if errors are found). This separation ensures that proof synthesis and evaluation remain distinct while still enabling iterative refinement. Finally, it decides when to stop the refinement loop. Termination occurs either when the Verifier certifies the proof as complete and error‐free, or when the number of attempts exceeds a configurable threshold.
@@ 3.1.2 Prover
The Prover is the constructive core of the system. Its task is to transform unproven Lean theorems into completed proofs. Theorem proving requires both creativity – finding the right lemma or using the right tactic – and rigor – ensuring that the structure and Lean code are syntactically correct. To achieve this, the Prover balances LLM-based heuristic exploration with rigorous Lean formalization aided by the MCP Lean tools made available by the lean-lsp-mcp (see Section 3.2).
We instruct the Prover to carry out its task following an incremental, step-by-step approach, and to write each update to the theorem proof to a .lean file. This is for two reasons: first, it satisfies the requirements of the MCP Lean tools, some of which need a .lean file path to inspect the code contained within; second, it allows the user to inspect the proof process in real time. In Figure 2, we present Lean code snippets at important stages of this process. Initially, the Prover identifies target theorems, by scanning Lean files for unfinished proofs marked with sorry, a placeholder tactic indicating an incomplete proof (top-left). Then, it writes a proof sketch, a coarse-grained natural language outline of the proof’s logical flow which breaks down a complex proof into more manageable steps (top-right). Next is formalization, where each step in the sketch is formalized into a Lean statement starting with have and ending with sorry (bottom-left); this brings into Lean’s proof context new statements that are to be proven from earlier hypotheses of the theorem. Then, the Prover goes through each step sequentially, proposing Lean tactics to substitute each sorry. After completing each step, the Prover uses a specific Lean tool, lean_diagnostic_messages (see Section 3.2) to assess if the generated step is correct. If critical errors are detected or a sorry is still present, the Prover attempts to fix the error or correct its reasoning. When all the steps are correctly solved, the Prover ends its task (bottom-right).
Tool usage is crucial for the Prover to perform its task. This is clearly illustrated in Figure 1 (right), which is extracted from the LLM log during an experimental run. The figure shows how both exploration and formalization are achieved through tool-enhanced reasoning, where the Prover uses mcp tools to interact with Lean files (read_file and edit_file); identify goals at various points in the proof (lean_goal); search for theorems in Mathlib (lean_leansearch); and verify the correctness of its proofs (lean_diagnostic_messages) This approach allows the Prover to function like an automated yet cautious mathematician: it drafts a plan, incrementally explores and implements ideas, verifies their correctness in Lean, and advances only once each step has been validated.

### v3

@@ Abstract
We present Ax-Prover, a multi-agent system for automated theorem proving in Lean that can solve problems across diverse scientific domains and operate either autonomously or collaboratively with human experts. To achieve this, Ax-Prover approaches scientific problem solving through formal proof generation, a process that demands both creative reasoning and strict syntactic rigor. Ax-Prover meets this challenge by equipping Large Language Models (LLMs), which provide knowledge and reasoning, with custom Lean tools which ensure formal correctness. To evaluate its performance as an autonomous prover, we benchmark our approach against both frontier LLMs and specialized prover models on two public math benchmarks and on two Lean benchmarks that we introduce in abstract algebra and quantum theory. On the public datasets, Ax-Prover is the top prover among those that do not rely on any domain-specific training; on the new benchmarks, it largely outperforms all baselines. This shows that, unlike specialized systems that struggle to generalize, our tool-based agentic theorem prover approach offers a generalizable methodology for formal verification across diverse scientific domains. Furthermore, we showcase Ax-Prover as a researcher-friendly assistant through two practical case studies in classical and quantum cryptography, pillars of secure communication, where it worked with domain experts to formalize and verify challenging security guarantees using standard human-agent interactions, enabling those without expertise in Lean to engage in this burgeoning domain.
**footnotetext: Equal contribution, authors listed alphabetically. Benjamin initiated the project and led early development; including curating the AbstractAlgebra dataset; Marco implemented the AX-Prover and led the writing of the paper; Jacob ran most of the experiments and significantly contributed to the writing and ideas of the paper.
@@ 1 Introduction
Developing Large Language Models (LLMs) that can reason reliably across scientific domains remains a central challenge for AI, both in academia and in industry. LLM-based formal reasoning systems have mainly been developed for mathematics, where they have achieved outstanding results [chervonyi2025gold, chen2025seed]. Recently, considerable effort has been put into training reasoning LLMs for formal theorem proving using Lean [moura2015lean], an open-source programming language and interactive proof assistant. Together with its community-driven Mathlib library [mathlib-stats], Lean provides a rigorous setting where AI systems must engage with symbolic reasoning and structured formalization, building on an evolving body of mathematical knowledge. LLM provers such as the DeepSeek-Prover series [deepseek-prover, deepseek-prover-v1, deepseek-prover-v2], Kimina-Prover-72B [kimina-prover], Goedel-Prover [lin2025goedel, lin2025goedel2], and Seed-Prover [chen2025seed] have shown that specialized prover models can be distilled from frontier LLMs and trained for theorem proving in Lean, reaching state-of-the-art performance on math benchmarks like MiniF2F [minif2f] and PutnamBench [tsoukalas2024putnambench]. Despite these results, these models face some key limitations. First, since they were mainly trained and tested on the domain of mathematics, their ability to generalize beyond this domain remains unclear. Relatedly, they are usually trained on fixed versions of the fast-evolving Mathlib library, making them brittle to changes introduced in new Mathlib versions, such as new or renamed definitions. Keeping them up to date would require frequent re-training and systematic “forgetting” of out-of-date knowledge, which adds significant cost. [^1] For example, Deepseek-Prover-V2-671B was released on April 30, 2025 and, in our experiments, we observed that it uses the lemma sqrt_eq_iff_sq_eq, which was deprecated in favor of sqrt_eq_iff_eq_sq on March 3, 2025. Second, while training improves their ability to produce Lean proofs, it narrows their capabilities relative to general purpose LLMs as they become unable to use external tools and engage in human-AI collaboration. Finally, it is hard to run them as they require high-performance computing and expertise to be successfully deployed and used. Together, these issues suggest that scaling increasingly large specialized provers may yield diminishing returns in both flexibility and usability.
@@
In contrast, general-purpose LLMs like Claude [claude_docs] and GPT [openai_models_docs] encode substantial knowledge across diverse domains (e.g., mathematics, physics, and computer science), while also exhibiting strong natural language understanding, problem solving skills, and interaction capabilities. Moreover, they are easily accessible through APIs, making them convenient to deploy and integrate into any workflow. Yet, they are not explicitly trained to formalize statements or construct proofs in Lean, nor can they natively interact with the Lean environment. This creates a sharp division: specialized provers are tightly integrated with Lean but narrow in scope and hard to use, whereas general-purpose LLMs are broad in scope and easily accessible but lack the ability to interface with the formal reasoning infrastructure required for theorem proving.
@@
To address this gap, we introduce Ax-Prover,222“Ax” stands for “axiomatic”, reflecting the base principles in mathematics and physics, the domains explored in this work. a new agentic workflow for theorem proving in Lean that leverages the Model Context Protocol (MCP) [modelcontextprotocol2024] to equip general-purpose LLMs with Lean tools from the lean-lsp-mcp repository [lean-mcp]. Ax-Prover combines the reasoning capabilities of LLMs with the formal verification power of Lean. The LLM analyzes unproven theorems, proposes proof sketches, and generates step-by-step Lean code, while the Lean tools allow the LLM to inspect goals, search for relevant results, locate errors, and verify proofs – capabilities essential for rigorous formal theorem proving.
@@
Ax-Prover overcomes the main limitations of state-of-the-art provers. First, using frontier LLMs prevents domain overspecialization while the MCP interface allows the system to work with any recent version of Mathlib and additional library dependencies including custom definitions relevant to the project without needing to be retrained. Second, it preserves tool-use and conversational abilities, enabling interactive collaboration with humans. Third, by leveraging existing frontier models, it sidesteps the need to host or deploy specialized systems.
@@
We evaluated Ax-Prover on two public datasets of mathematics competition problems (NuminaMath-LEAN [numinamath-lean] and PutnamBench [tsoukalas2024putnambench]) and introduce two new datasets to enable evaluation in new domains. The first of these, AbstractAlgebra, focuses on algebraic structures such as groups, rings, and fields, testing the provers’ abilities to reason in a more abstract, research-oriented setting rather than the competition-driven style of existing datasets. The second new dataset is QuantumTheorems, which represents an initial step toward automated theorem proving in a scientific domain beyond pure mathematics, evaluating the models’ formal reasoning in quantum mechanics. Our results show that Ax-Prover has competitive performance on PutnamBench – achieving the highest accuracy for fully open-sourced agents [^3] We use “open-sourced” according to the usage in the official leader-board [putnambench_leaderboard]. – and outperforms general-purpose LLMs not equipped with Lean tools as well as state-of-the-art specialized provers on the other datasets, with a significant margin on the ones we introduce.
Besides functioning as an autonomous solver, Ax-Prover was also designed to serve as an assistant for human researchers. To demonstrate these capabilities, we present two researcher-facing use cases in the field of cryptography in Sections LABEL:sec:Cryp and LABEL:sec:QKD. Cryptography is an ideal proving ground for Lean because its security guarantees depend on precise mathematical reasoning, yet the field often lacks standardized assumptions and explicit logical structure. Machine-checked proofs can therefore transform the way such guarantees are built and trusted – ensuring that every step, assumption, and reduction is explicit and verifiable. In the first use case, Ax-Prover collaborates with a cryptography researcher to formalize and verify an alternate definition for the branch number of a matrix [mishra2024new], revealing a subtle gap in the informal argument and producing a reusable Lean certificate on their own laptop in 2 days’ time. In the second, it assists quantum-information researchers in formalizing an entropy bound in quantum key distribution [lo1999unconditional], converting a physics-style derivation into a machine-checked component. Together, these examples show that beyond benchmark accuracy, Ax-Prover lowers the practical barrier for researchers to use Lean in their own work, brings clarity and rigor to complex reasoning, and enables interpretable, researcher-driven verification for security-critical domains.
@@
Our contributions are threefold: (i) We design Ax-Prover, a lightweight agentic workflow that connects general-purpose LLMs to Lean tools via MCP, and demonstrate that it performs on par with or surpasses both general-purpose LLMs and specialized provers across several scientific domains; (ii) We introduce new formalized Lean datasets covering abstract algebra and quantum physics, complementing existing benchmarks; (iii) We showcase Ax-Prover’s capabilities as an assistant through use cases where the system collaborated with domain experts to formally verify the proof of a recent cryptography result [mishra2024new] as well as an entropy bound for the Lo-Chau security framework for Quantum Key Distribution (QKD) [lo1999unconditional].
@@ 2 Related Work
Automated theorem proving in Lean has roots in classical approaches such as decision procedures [demoura2008z3, barbosa2022smt] and heuristic-guided proof search [kovacs2013vampire, schulz2019e]. However, these approaches face particular challenges: the former does not handle general mathematical domains (e.g. transcendental functions and complex numbers) and the latter does not perform well out of distribution. More recent work integrates machine learning: from heuristic tuning [urban2011machine] to premise selection and tactic prediction [irving2016deepmath, huang2019holstep], culminating in transformer-based language models capable of generating Lean proofs [polu2020generative, lample2022deep, polu2023formal, xin2024lean]. More recent large-scale systems extend this trend by training LLMs on formal proving through distillation, supervised finetuning, and reinforcement learning. Current examples of specialized models are Kimina-Prover [kimina-prover], the DeepSeek-Prover family [deepseek-prover, deepseek-prover-v1, deepseek-prover-v2], Goedel-Prover 1 and 2 [lin2025goedel, lin2025goedel2], Prover Agent [baba2025prover], Apollo [ospanov2025apollo], and Seed-Prover [chen2025seed]. All of these are highly specialized provers, which take in a Lean theorem as input and autonomously produce a proof. A very recent line of work is exploring agentic flows that include frontier LLMs and a formal verifier; examples include Hilbert [varambally2025hilbert] and Aristotle [achim2025aristotle]. Although we also adopt a similar approach, some crucial differences exist, namely: (i) we give the LLM direct access to Lean tools via MCP; (ii) our framework requires neither training nor finetuning [achim2025aristotle] and does not rely on any specialized provers [varambally2025hilbert]; (iii) we demonstrate the effectiveness of our approach on domains beyond just mathematics; (iv) we showcase the capabilities of our system as an interactive assistant to human researchers.
Finally, a parallel line of work has explored classical machine learning for supporting experts in theorem proving in Lean, for example, in premise selection and tactic prediction [gauthier2021tactictoe, blaauwbroek2020tactician], and more recently through LLMs that connect to Lean via external interfaces [ayers2023leanlm, azerbayev2023llm, song2024lean]. These approaches illustrate the promise of AI-assisted proving, but they remain resource-intensive and difficult to adapt across scientific domains. Recent efforts, such as [kumarappan2024leanagent], attempt to remedy this by emphasizing greater adaptability within Lean. At the same time, there is growing interest in human-AI collaboration: conversational assistants [collins2023llm] and “copilot”-style integrations [chen2021copilot] suggest how formal tools can augment, rather than replace, human reasoning. Our work builds on this trajectory by closing the gap between heavyweight, specialized provers and lightweight, researcher-friendly systems that can be more readily adapted to the evolving Lean ecosystem.
@@ 3 System Architecture
We implement Ax-Prover as a multi-agent architecture with three agents, each of which is implemented as an LLM with specific prompts: the Orchestrator, the Prover, and the Verifier. Following recent agentic designs for complex tasks such as scientific discovery [gottweis2025towards, yamada2025ai], we avoid a monolithic setup by assigning each specialized agent a distinct role. This separation enables specialization and modularity: agents can be independently optimized, replaced, or extended, allowing researchers to adapt Ax-Prover to their own needs without destabilizing the system.
@@ 3.1.1 Orchestrator
The Orchestrator’s role is analogous to a scheduler in distributed systems: it does not perform computation itself but ensures that computation flows smoothly across agents. It holds three main responsibilities. First, it handles task assignment, as it receives user input and instructs the Prover accordingly. Next, it manages feedback routing by taking diagnostic outputs from the Verifier and giving structured feedback to the Prover (if errors are found). This separation ensures that proof synthesis and evaluation remain distinct while still enabling iterative refinement. Finally, it decides when to stop the refinement loop. Termination occurs either when the Verifier certifies the proof as complete and error-free, or when the number of attempts exceeds a configurable threshold.
@@ 3.1.2 Prover
The Prover is the constructive core of the system. Its task is to transform unproven Lean theorems into completed proofs. Theorem proving requires both creativity – finding the right lemma or using the right tactic – and rigor – ensuring that the structure and Lean code are syntactically correct. To achieve this the Prover balances LLM-based heuristic exploration with rigorous Lean formalization aided by the MCP Lean tools made available by the lean-lsp-mcp (see Section 3.2).
We instruct the Prover to carry out its task following an incremental, step-by-step approach, and to write each update to the theorem proof to a .lean file. This is for two reasons: first, it satisfies the requirements of the MCP Lean tools, some of which need a .lean file path to inspect the code contained within; second, it allows the user to inspect the proof process in real time. In Figure 2, we present the main stages of the Prover process. Initially, the Prover identifies target theorems by scanning the input Lean file for unfinished proofs marked with sorry, a placeholder tactic indicating an incomplete proof (top-left). Then, it writes a proof sketch, a coarse-grained natural language outline of the proof’s logical flow which breaks down a complex proof into more manageable steps (top-right). Next is formalization, where each step in the sketch is formalized into a Lean statement starting with have and ending with sorry (bottom-left); allowing the prover to see the original theorem statement broken down into current and next steps within the lean context. Then, the Prover goes through each step sequentially, proposing Lean tactics to substitute each sorry. After completing each step, the Prover uses a specific Lean tool, lean_diagnostic_messages (see Section 3.2) to assess if the generated step is correct. If critical errors are detected or a sorry is still present, the Prover attempts to fix the error or correct its reasoning. When all steps are correctly solved, the Prover ends its task (bottom-right).
Tool usage is crucial for the Prover to perform its task. This is clearly illustrated in Figure 1 (right), which is extracted from the LLM log during an experimental run. The figure shows how both exploration and formalization are achieved through tool-enhanced reasoning, where the Prover uses MCP tools to interact with Lean files (read_file and edit_file); identify goals at various points in the proof (lean_goal); search for theorems in Mathlib (lean_leansearch); and verify the correctness of its proofs (lean_diagnostic_messages). This approach allows the Prover to function like a cautious mathematician: it drafts a plan, incrementally explores and implements ideas using relevant tools, verifies their correctness in Lean, and advances only once each step has been validated.